* LATEST HEMP NEWS *
Last update: Autumn 2004


2004 September

Summary of HEMP Activities as of Summer 2004
A summary of the HEMP major activities , major findings , and publications are available. Note that these cover the initial activities of HEMP - there is yet much to do and learn. At this time, HEMP is on temporary hold.


2004 Spring

Moving pains...
A large number of the "Latest News" items were lost during the transition from Univerity of Texas to San Diego State University (an unexpected loss of the host computer, and no backups of this Latest News directory). I will not attempt to reconstruct the news items from memory. News items lost span 2000 April through 2003. Apologies to all involved.


2000 Apr 17

HEMP joins the Akn 564 Campaign
HEMP has officially joined the International AGN Watch Campaign on Akn 564. Akn is a Narrow Line Seyfert 1 AGN and will be observed this spring/summer with HST, RXTE, Chandra, ASCA and XMM. HET time has been officially awarded.
The successful "trial run" on Mrk 817 will continue, but at a lower priority. To date, we have obtained 18 spectra since Feb 7, giving an average of one spectra per 3.8 days. We will keep this program running through the summer.


2000 Mar 17

QSOM update
The QSO Monitoring part of HEMP is rolling along. We have runs at McDonald scheduled for April, May and June. We have already had 2 good runs this year. On the down side, the queue-scheduled 0.9m telescope SSTO program at CTIO has been cancelled by NOAO, so there will be no more data from CTIO after this semester. On the bright side, we are obtaining data from St. Andrews University (Scotland) to support our Mrk 817 campaign. Keith Horne formally joined HEMP in February, and his student Rachel Street has been obtaining CCD images for us with the JGT telescope. Here at UT, an undergrad, Margaret Moerchen, continues to plug away at the many, many images we've obtained; she will be taking the April run at McDonald.
As for the HET, there still appear to be some throughput problems (perhaps due to bad tracking/guiding?). Luckily, the Seyferts we are observing this year are so bright (~15th mag) that this is not at all an issue.


2000 Mar 13

"Baby's first steps..."
After some nasty trouble with the guide camera in February, the HET+LRS are back up and running. To date, we have 8 spectra (about one spectrum every 4.5 days on average). A modest start, but even a journey of a thousand miles starts with one step. While the spectra produced in the Fall of 1999 could be regarded as those that "only a mother could love", these Mrk 817 data are superb.
It is becoming clear that it will soon be quite a chore to keep up with the data flow rate. Unfortunately, a simple "pipeline" data reduction process does not appear possible.
On another front, we have had 2 highly successful QSO Monitoring runs this year at McDonald, with 3 more planned in the Spring. Also, data from the 0.9mQ SSTO program at CTIO are now coming in as well.


2000 Feb 15

The first HEMP spectra have been obtained!
The HEMP is in the queue and data are being obtained. If the weather cooperates, we expect a spectrum every 2 days. The data look great - a S/N of over 100 in the continuum has been measured in a 9 minute exposure.


2000 Jan 20

HEMP is #1!
The U.T. HET TAC has given the HEMP proposal its highest (non-T.O.O.) ranking! The HEMP proposal for Feb-Apr has been given top-priority time, so we will begin our very first campaign in February. HEMP is alive and kicking!


2000 Jan 06
We are very happy to report that the HET+LRS are now in routine operation!!

In fact, the HET+LRS have been taking data since 1999 October, and the queue mode of observing is working wonderfully. For more on the HET itself (general info, instruments, observing schedule, etc.) look at the Hobby-Eberly Telescope page; for more specific info on the current instrumentation, see the LRS, UFOE & HRS page.

We have obtained a considerable amount of (non-HEMP) data and have been busy shaking out the bugs and calibrating these spectra. While things are not perfect, the data we have already obtained show that the HET and LRS are in a state that is completely adequate for the HEMP.

Currenly, the HET is operating only during dark and grey time, with engineering shut-down occuring during bright time. This is of course a problem for echo mapping since we need a well-sampled time series. However, the required engineering time is decreasing each month (and we can use other telescopes at McDonald to suppliment the HET observations if necessary). So we expect a highly productive year for the HEMP. Indeed, we will begin a small-scale echo mapping project this spring, and the full-blown Project will begin this summer.

This figure shows a LRS spectrum of the Seyfert 1 galaxy Mrk 110. This bright AGN (V ~15th mag) is very similar to the kind of low-redshift AGN we will echo map. The S/N of this spectrum is >100 in this 10 min exposure. (Data courtesy of W. Kollatschny).

Another view of Mrk 100, this time showing the region near Hbeta. The Hbeta line is relatively weak and narrow in this spectrum. For HEMP targets, we expect the line to be broader and stronger. This spectrum also shows the strong, non-variable forbidden [OIII] lines we will use as internal flux calibrators. The HeII 4686 line is present, which we will also attempt to map. For this AGN, our HET observations show considerable line profile changes in both Hbeta and HeII over the span of 3 months.


QSO Monitoring Project Status (QSOM):
The high-z QSO Monitoring is moving right along.
(1) We have been awarded time in the winter/spring at CTIO and at McDonald for the QSOM.
(2) We are currently reducing the data from last spring. The initial calibrations show that we can achieve better than a few percent photometry on a V=17th mag QSOs even during grey/bright time. Click here for a sample light curve The data reduction has been carried out by an undergrad at UT, M. Moerchen. For a description of her work, click here.


Back to the HEMP home page.